# Diverging Towards Hallucination: Detection of Failures in Vision-Language Models via Multi-token Aggregation

### v1

@@ 7 Limitations
MTRE does have some limitations. First, it requires white-box access to the full sequence of early logits, so it cannot be applied when only final outputs or API-level confidences are available. Second, all experiments are conducted on four 7B-parameter open-source VLMs and a handful of vision-question benchmarks; MTRE’s generality to much larger models, other modalities (e.g. video), non-English prompts, or truly "in-the-wild" user queries remains untested. Third, like other white-box probes, MTRE can be sensitive to prompt phrasing and decoding strategies (greedy vs. top-$p$), so its robustness under different prompting schemes warrants further study.
@@ 8 Conclusion
Despite the impressive capabilities of vision-language models (VLMs), their vulnerability to hallucinated or unsafe outputs continues to hinder their reliability in real-world, safety-critical settings. These risks are especially pronounced when models encounter ambiguous or adversarial inputs, where traditional output-level assessments may fail to detect potential failures. Addressing this challenge requires methods that go beyond surface-level outputs to better capture the underlying reasoning dynamics of the model.
@@
More broadly, our work supports the growing recognition that true safety in AI systems must be grounded in a deep, introspective understanding of model behavior rather than superficial performance metrics alone. By demonstrating that internal signals—specifically, token-level logits—carry actionable information about output reliability, we pave the way for more interpretable evaluation techniques that can curb hallucinations, reduce the spread of misinformation, and build user trust in applications as critical as healthcare and education. At the same time, it is important to remain vigilant: adversaries may learn to game these detectors, and overreliance on automated checks can foster complacency, allowing subtler failures to slip through when stakes are highest. Striking the right balance between automated monitoring and human oversight will be essential to align ever more powerful models with human values and deploy them responsibly across diverse real-world contexts.
Acknowledgments
This manuscript has been assigned LA-UR-25-24783. This research was funded by the Los Alamos National Laboratory (LANL) Laboratory Directed Research and Development (LDRD) program under grants 20230287ER, 20250850ECR and 20240868PRD3 and supported by LANL’s Institutional Computing Program, and by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. We also thank Sambanova Systems for providing access to API calls for dataset synthesis utilized in this work.
@@ Appendix A Datasets
We primarily evaluate or improvements on the datasets utilized by a first token linear probing technique discussed in Zhao et al. [2025]. For each dataset, we construct a separate Type 2 dataset in the main text.
MM-SafetyBench
MM-SafetyBench applies jailbreaking attacks to LVLMs across thirteen scenarios using malicious text prompts and images Liu et al. [2023a]. The original dataset includes 1,680 unsafe questions for attacks, with each question generating three types of images: one created by Stable Diffusion Podell et al. [2023], one with rendered text, and one combining the first two. For our work, we use the augmented version of this dataset introduced in Zhao et al. [2025], which balances the dataset by adding a new data generation pipeline in MM-SafetyBench. This pipeline generates a total of 1,800 safe question-image pairs through GPT-4 API/ Llama3-70B deployed on SambaNova cloud API prompts covering topics such as daily activities, economics, physical health, legal matters, politics, finance, sex, and government.
We train all models (MMD, Linear Probing, and MTRE) on these data to distinguish whether the output will be harmful. To remain consistent with Zhao et al. [2025], we also randomly select 10 samples from each category in both safe and unsafe sets and use 90 safe and 130 unsafe samples for training. The remaining data of the augmented MM-SafetyBench is used as the test set.
@@
MAD-Bench consists of 850 image-question pairs designed to deceive LVLMs. These deceptive pairs target various aspects, including object count, non-existent objects, object attributes, scene understanding, spatial relationships, and visual confusion Li et al. [2023]. For example, given an image of two cats, a deceptive question might be: ’What are the three cats doing?’ In this case, rather than answering the question directly, the model should recognize the inconsistency between the question and the image. We also utilize an augmented dataset which adds an additional generated 1,000 normal questions based on the COCO val2017 dataset. We use 100 deceptive and 100 normal samples to train each proposed technique. The remaining data is then used as a validation dataset in each of our experiments.
@@
We use a third dataset, the MathVista dataset Lu et al. [2023], which contains 1,000 image-question pairs related to math problems. This dataset challenges the model by requiring it to predict various types of answers, such as multiple-choice options, floating-point numbers, integers, and lists, making correctness prediction more complex. We prompt VLMs with the math visual prompts and evaluate their accuracy using GPT-4, following the scripts provided in the official GitHub repository.
Given the limited size of the dataset we implement a 4-fold cross-validation method (in contrast to the 10-fold selection in Zhao et al. [2025]) to ensure the robustness of our analysis. In each fold, the model is provided with the output logits and trained to predict the accuracy of responses based on the logit distribution of each output token. Once trained, the model is applied to predict the accuracy of responses in the test segment. The performance of the model on this dataset is evaluated using the metrics discussed in Section 6.2 across all folds.
@@
Below we note the descriptions of the datasets given by Rahmanzadehgervi et al. [2024]. Note that we alter each dataset primarily to experiment with more data, and more complicated cross-validation splits. We reduced the amount of shapes/diversity in all shape datasets due to the difficulty for smaller open-source models, and to reduce the mode collapse in VLM predictions. Similar to MathVista we implement a 4-fold cross-validation to account for the size of dataset. We are careful to not make each of the training splits identical to any of the validation splits for any of the folds.
  • Intersecting Lines: Following the work of Rahmanzadehgervi et al. [2024] we create 600 images of 2D line plots drawn on a white canvas. Each line plot consists of two line segments, defined by three points whose x-coordinates are fixed and equally spaced. The y-coordinates are randomly sampled to create two plots that intersect at exactly 0, 1 or 2 points. The goal of the VLM is to count the number of line intersections. There are 200 images with 0 intersections, 200 with 1 intersection, and 200 with 2 intersections. We denote explicit configurations in practice below:
@@
    • Shape Placement: The center of the nested stack is randomly positioned within the range $[-5,5]$ for both $x$ and $y$ coordinates.
@@
Appendix B Model specific details
B.1 Training Protocol
[…]
B.2 Considerations for uneven sentences
B.3 Masking tokens
Given that the length of sentences produced by VLMs may vary wildly, we experiment with at most 10 output tokens. In practice, sentences shorter than 10 tokens require zero padding for missing logits. Therefore, we begin by defining an $\epsilon$-norm mask $m_{\ell}=\mathbf{1}[\lVert x_{\ell}\rVert_{2}>\epsilon]\!.$. Below we redefine section LABEL:sec:method, to improve reproducibility. For every prefix length $k\!\in\!\{1,\dots,K\}$ (with $K\!\leq\!10$ in experiments) we compute the masked log likelihood under each hypothesis:
$\ell^{(k)}_{1}$ $=\sum_{\ell=1}^{k}m_{\ell}\,\log p_{\ell},$ $\ell^{(k)}_{0}$ $=\sum_{\ell=1}^{k}m_{\ell}\,\log(1-p_{\ell}).$ (6)
Their difference is the cumulative log-likelihood ratio (LLR):
$\boxed{\,\Lambda^{(k)}=\ell^{(k)}_{1}-\ell^{(k)}_{0}=\sum_{\ell=1}^{k}m_{\ell}\log\!\frac{p_{\ell}}{1-p_{\ell}}\,}.$
Equation (1) is computed in the validation loop by summing $\log p_{\ell}$ and $\log(1-p_{\ell})$ across locations and subtracting the two running totals. We hypothesize that tuning early stopping on certain tokens may be an option to improve performance further (i.e stop the accumulation before K), but we leave the investigation and validation of early stopping for future work.
B.4 Hyperparameters
All experiments for Arithmetic tasks as discussed in Table 7 can be reproduced using the following hyper parameters: $input\_dim=32000$, $embed\_dim=512$, $num\_heads=8$, $num\_layers=3$, dropout$=0.1$, epochs$=100\text{ to }300$, $batch\_size=32$, lr$=1\times 10^{-5}$, $token\_level=10$.
We utilize Binary cross entropy loss and Adam for our optimizer. Hyper-parameters for other datasets will be officially released in supplemental material.
Appendix C Results
Table 5: Comparative Performance Metrics Across Vision-Language Models for final classification performance on the original benchmark using Type 1 responses.
Model Method Safety MAD
Acc Auc F1 Acc Auc F1
OE
LLAVA-7B Linear Probing 90.55 97.31 90.71 92.22 97.16 92.16
P(True) 57.30 56.45 67.03 73.17 80.81 73.24
MTRE 90.95 96.90 90.91 92.06 97.33 91.92
LLAMA-Adapter Linear Probing 90.31 95.99 90.80 91.28 97.10 91.08
P(True) 67.33 72.13 69.80 54.78 57.42 63.07
MTRE 90.25 96.73 90.64 90.37 96.36 90.79
MPLUG-Owl Linear Probing 92.24 97.58 92.60 94.28 98.74 94.21
P(True) 73.04 81.08 75.13 76.89 85.41 74.91
MTRE 91.29 96.98 89.31 93.72 98.13 92.46
MiniGPT4 Linear Probing 86.66 94.86 87.11 84.50 91.73 84.37
P(True) 55.06 51.16 68.88 49.17 49.36 50.35
MTRE 86.17 94.04 86.19 81.67 88.85 80.85
OEH
LLAVA-7B Linear Probing 87.67 95.46 87.82 90.00 95.61 89.88
P(True) 52.09 49.97 67.71 73.28 79.31 76.66
MTRE 88.19 95.42 88.35 87.44 94.98 87.10
LLAMA-Adapter Linear Probing 88.96 95.34 89.30 87.89 95.05 87.50
P(True) 65.06 72.09 58.60 62.36 72.17 65.61
MTRE 89.11 95.66 89.60 86.11 93.84 86.00
MPLUG-Owl Linear Probing 89.94 96.73 90.24 93.33 98.14 93.25
P(True) 77.49 78.54 76.38 73.83 76.34 76.65
MTRE 90.46 96.63 90.73 91.61 97.41 91.67
MiniGPT4 Linear Probing 83.22 92.03 82.99 86.94 93.72 86.62
P(True) 54.60 58.82 52.75 54.06 53.89 31.37
MTRE 85.52 93.00 86.42 81.28 88.74 81.33
MQ
LLAVA-7B Linear Probing 91.01 96.97 91.40 90.83 96.86 90.78
P(True) 49.85 52.33 57.72 55.17 58.09 60.38
MTRE 90.95 97.01 91.36 91.78 97.26 91.69
LLAMA-Adapter Linear Probing 88.77 95.51 88.98 87.44 95.32 87.20
P(True) 60.52 65.01 52.53 54.31 54.21 45.72
MTRE 90.74 97.24 91.20 88.50 95.50 88.22
MPLUG-Owl Linear Probing 91.99 97.65 92.24 93.66 98.40 93.54
P(True) 65.12 71.39 62.71 83.33 89.19 84.65
MTRE 91.07 97.21 91.40 92.28 96.46 93.08
MiniGPT4 Linear Probing 81.93 90.03 82.09 83.00 90.52 82.47
P(True) 47.12 47.65 57.18 56.94 62.16 64.69
MTRE 82.88 91.29 82.96 82.28 90.00 82.42
Table 6: Comparative Performance Metrics Across Vision-Language Models for Self-Evaluation Type 2 responses.
Model Method Safety 2 MAD 2
Acc Auc F1 Acc Auc F1
OE
LLAVA-7B Linear Probing 48.65 49.20 57.90 64.55 55.73 76.28
P(True) 39.29 42.34 23.09 68.17 60.98 35.98
MTRE 68.58 58.83 78.64 75.56 63.18 85.48
LLAMA-Adapter Linear Probing 58.44 54.18 69.47 87.17 93.52 87.00
P(True) 35.80 40.80 24.08 51.61 53.72 64.02
MTRE 66.38 48.32 79.79 87.44 93.57 87.40
MPLUG-Owl Linear Probing 48.04 40.44 60.59 80.83 86.52 84.38
P(True) 71.06 63.55 82.91 67.17 65.38 71.92
MTRE 71.87 55.17 82.89 84.44 89.47 87.54
MiniGPT4 Linear Probing 64.29 66.29 71.17 76.38 78.51 73.97
P(True) 46.20 50.96 46.30 53.33 56.13 45.88
MTRE 69.39 74.85 76.64 77.78 83.40 77.06
OEH
LLAVA-7B Linear Probing 46.84 55.38 32.91 81.28 88.54 81.83
P(True) 42.36 36.67 14.55 51.67 57.78 61.23
MTRE 53.71 55.69 48.86 81.17 89.82 81.17
LLAMA-Adapter Linear Probing 45.21 46.62 43.23 87.56 94.58 87.40
P(True) 53.16 51.91 60.29 54.31 50.68 27.85
MTRE 65.49 64.19 64.59 86.00 91.49 87.01
MPLUG-Owl Linear Probing 44.29 49.54 38.48 60.11 70.93 74.97
P(True) 71.81 73.73 73.89 67.28 69.24 73.95
MTRE 68.53 67.64 70.05 73.83 82.16 79.17
MiniGPT4 Linear Probing 55.80 60.96 41.77 70.50 78.01 72.44
P(True) 65.61 66.94 59.22 49.94 53.41 63.97
MTRE 69.02 75.06 64.44 75.72 84.36 77.01
MQ
LLAVA-7B Linear Probing 43.68 40.72 43.96 85.88 93.19 86.75
P(True) 46.01 42.74 39.31 51.72 49.85 59.64
MTRE 57.39 57.66 60.42 86.28 94.47 86.91
LLAMA-Adapter Linear Probing 61.84 63.16 50.60 79.28 87.18 79.58
P(True) 45.49 50.52 56.39 54.05 53.82 57.14
MTRE 70.31 73.20 56.24 78.39 87.02 78.95
MPLUG-Owl Linear Probing 74.05 81.64 68.02 89.67 95.86 90.55
P(True) 64.23 71.35 59.99 80.33 84.01 83.67
MTRE 79.39 83.47 77.48 92.39 96.46 93.09
MiniGPT4 Linear Probing 56.71 59.12 53.17 65.50 69.42 68.81
P(True) 48.83 45.86 38.36 48.78 49.43 54.08
MTRE 66.26 67.77 57.85 72.59 73.83 72.59
Model Method Circle Triangle Lines Math Squares
Acc AUC F1 Acc AUC F1 Acc AUC F1 Acc AUC F1 Acc AUC F1
LLAVA-7B LR 73.73 79.00 75.68 77.73 74.13 83.94 70.50 69.20 36.97 73.30 71.68 38.49 61.20 59.77 49.18
$\pm$11.25 $\pm$13.53 $\pm$18.05 $\pm$7.12 $\pm$24.68 $\pm$4.37 $\pm$6.35 $\pm$6.00 $\pm$12.19 $\pm$1.24 $\pm$1.73 $\pm$4.14 $\pm$22.56 $\pm$34.91 $\pm$15.95
P(True) 81.09 77.00 85.97 91.46 80.79 94.31 48.83 54.92 39.85 71.00 63.14 14.41 60.50 64.77 64.83
$\pm$14.08 $\pm$4.27 $\pm$1.95 $\pm$14.18 $\pm$32.90 $\pm$9.38 $\pm$8.81 $\pm$12.72 $\pm$3.54 $\pm$0.83 $\pm$1.49 $\pm$5.80 $\pm$6.40 $\pm$8.71 $\pm$7.48
MTRE 82.00 86.73 82.73 87.45 89.10 91.28 75.33 71.21 36.55 75.70 74.48 35.69 87.50 94.15 75.84
$\pm$13.63 $\pm$13.15 $\pm$16.17 $\pm$15.11 $\pm$21.80 $\pm$6.92 $\pm$4.69 $\pm$12.13 $\pm$16.40 $\pm$1.73 $\pm$2.31 $\pm$9.64 $\pm$11.06 $\pm$6.04 $\pm$27.91
LLAMA-Adapter LR 85.82 93.30 86.18 88.82 87.89 91.73 90.80 97.83 85.09 72.80 67.84 35.50 71.83 37.88 25.41
$\pm$12.62 $\pm$8.54 $\pm$12.22 $\pm$8.13 $\pm$15.78 $\pm$7.21 $\pm$2.81 $\pm$1.94 $\pm$7.28 $\pm$2.17 $\pm$5.54 $\pm$5.08 $\pm$15.54 $\pm$22.01 $\pm$25.71
P(True) 58.46 71.89 66.13 68.73 59.00 60.69 59.00 50.66 28.52 69.00 68.52 43.78 42.50 56.08 41.85
$\pm$10.36 $\pm$17.86 $\pm$7.82 $\pm$2.78 $\pm$8.85 $\pm$7.60 $\pm$3.45 $\pm$7.84 $\pm$8.47 $\pm$1.08 $\pm$1.88 $\pm$2.11 $\pm$3.21 $\pm$4.45 $\pm$4.89
MTRE 93.27 97.98 93.26 94.45 93.61 96.40 93.50 97.59 89.92 75.30 67.57 35.87 89.33 87.16 82.36
$\pm$7.63 $\pm$3.50 $\pm$8.17 $\pm$9.38 $\pm$12.78 $\pm$5.90 $\pm$0.99 $\pm$1.34 $\pm$1.51 $\pm$2.96 $\pm$5.60 $\pm$8.77 $\pm$10.47 $\pm$13.20 $\pm$15.18
MPLUG-Owl LR 91.09 99.62 90.39 87.35 87.28 65.58 87.33 93.87 78.88 72.90 73.77 47.88 79.17 70.19 47.81
$\pm$6.58 $\pm$0.66 $\pm$90.39 $\pm$11.24 $\pm$17.66 $\pm$39.95 $\pm$3.40 $\pm$2.75 $\pm$9.81 $\pm$1.34 $\pm$2.93 $\pm$3.14 $\pm$10.29 $\pm$22.70 $\pm$24.96
P(True) 51.18 66.48 28.37 81.91 83.67 70.30 63.83 68.59 54.48 71.20 51.21 22.84 66.83 55.46 5.35
$\pm$10.37 $\pm$13.87 $\pm$28.37 $\pm$7.72 $\pm$13.03 $\pm$16.29 $\pm$4.86 $\pm$6.36 $\pm$1.47 $\pm$3.86 $\pm$3.86 $\pm$4.83 $\pm$31.26 $\pm$19.45 $\pm$3.28
MTRE 97.73 97.05 96.66 90.55 92.80 85.19 86.67 93.15 79.27 74.80 75.76 37.40 89.00 91.87 81.91
$\pm$3.94 $\pm$5.11 $\pm$5.79 $\pm$8.07 $\pm$8.77 $\pm$12.14 $\pm$3.37 $\pm$3.48 $\pm$4.77 $\pm$1.62 $\pm$2.79 $\pm$13.72 $\pm$7.00 $\pm$7.73 $\pm$9.72
MiniGPT4 LR 76.18 76.31 59.26 86.81 88.83 87.81 87.66 94.31 79.12 75.00 72.41 42.08 45.16 42.84 49.19
$\pm$8.23 $\pm$12.95 $\pm$23.64 $\pm$21.18 $\pm$17.93 $\pm$19.79 $\pm$3.90 $\pm$0.54 $\pm$3.41 $\pm$1.15 $\pm$2.52 $\pm$2.40 $\pm$2.72 $\pm$3.01 $\pm$6.71
P(True) 56.36 60.14 54.29 63.82 56.70 76.52 55.00 56.61 43.74 67.80 40.01 13.88 49.33 51.16 52.36
$\pm$9.25 $\pm$9.33 $\pm$13.26 $\pm$4.52 $\pm$2.19 $\pm$5.38 $\pm$2.60 $\pm$4.22 $\pm$3.26 $\pm$4.77 $\pm$3.64 $\pm$1.12 $\pm$4.32 $\pm$6.39 $\pm$5.57
MTRE 81.10 87.98 70.56 91.00 89.80 91.32 90.83 96.24 83.89 75.04 67.42 37.90 89.50 93.19 91.00
$\pm$17.31 $\pm$12.51 $\pm$31.57 $\pm$13.96 $\pm$15.21 $\pm$13.84 $\pm$1.19 $\pm$0.78 $\pm$2.49 $\pm$0.60 $\pm$2.59 $\pm$1.57 $\pm$7.34 $\pm$9.00 $\pm$5.42
Table 7: Performance comparison of methods using logits from multimodal models on different geometric shapes datasets, evaluated using Accuracy (Acc), Area Under the Curve (AUC), and F1 score. Reported values are means and standard deviations across cross-validation folds.
Table 8: Comparative Performance Metrics Across Vision-Language Models for Hallucination Detection on Type 1 responses.
Model Method Safety Hallu MAD Hallu
Acc Auc F1 Acc Auc F1
OE
LLAVA-7B Linear Probing 79.90 85.63 85.43 87.88 92.36 88.40
P(True) 31.14 44.29 00.00 72.67 79.30 75.74
MTRE 81.90 87.53 86.50 85.61 91.54 86.27
LLAMA-Adapter Linear Probing 79.78 85.73 85.00 90.27 95.88 90.18
P(True) 35.74 34.21 10.89 55.22 57.79 63.56
MTRE 81.10 87.12 86.67 85.27 85.25 92.64
MPLUG-Owl Linear Probing 83.65 89.20 88.66 88.0 93.26 88.72
P(True) 42.76 56.96 37.51 72.56 79.41 71.58
MTRE 84.11 88.45 89.55 85.11 90.99 85.74
MiniGPT4 Linear Probing 77.69 84.40 82.25 78.83 85.75 80.04
P(True) 62.27 51.62 75.46 49.00 49.14 51.12
MTRE 75.0 80.49 80.33 77.66 85.60 78.36
OEH
LLAVA-7B Linear Probing 71.84 70.34 80.88 72.55 69.04 81.74
P(True) 66.56 55.10 78.79 40.94 54.54 44.14
MTRE 82.82 86.60 88.67 82.38 85.73 88.65
LLAMA-Adapter Linear Probing 85.55 66.49 92.05 87.77 95.03 87.38
P(True) 50.68 57.88 62.71 62.36 72.17 65.61
MTRE 88.40 84.77 93.79 84.50 84.26 92.20
MPLUG-Owl Linear Probing 74.47 63.29 83.55 89.77 95.65 91.87
P(True) 51.69 67.20 52.94 59.83 54.86 67.21
MTRE 79.87 67.96 87.60 86.72 93.36 89.85
MiniGPT4 Linear Probing 64.90 61.17 76.40 86.77 93.68 86.46
P(True) 44.11 49.53 45.02 48.61 46.00 54.55
MTRE 71.84 70.87 81.37 80.11 88.07 79.58
MQ
LLAVA-7B Linear Probing 64.57 68.62 61.22 87.50 93.98 89.24
P(True) 50.03 48.95 37.42 58.17 58.67 65.76
MTRE 72.48 82.80 65.53 89.72 96.03 91.16
LLAMA-Adapter Linear Probing 88.71 95.52 88.53 79.77 46.13 88.04
P(True) 60.52 65.01 52.53 53.35 54.05 46.26
MTRE 91.28 97.30 90.82 78.38 81.44 87.11
MPLUG-Owl Linear Probing 86.87 92.95 88.51 88.00 93.26 88.72
P(True) 64.36 69.69 67.76 37.94 44.29 44.28
MTRE 90.39 95.41 91.65 85.11 90.99 85.74
MiniGPT4 Linear Probing 82.82 90.57 82.68 69.61 74.44 72.27
P(True) 47.12 47.65 57.18 48.50 54.25 30.67
MTRE 83.43 91.65 83.43 72.38 76.25 76.11
Table 9: We briefly investigate possible generalization improvements over using the first token through the arithmetic dataset. We monitor the auc score improvements for all arithmetic tasks. Absolute improvement in AUROC over single-token linear probing (Ours-LR). Gains are shaded green, losses red.
(a) Overlapping Circles & Triangles
Model Overlapping Circles Overlapping Triangles LLAVA-7B Adapter mPLUG MiniGPT4 LLAVA-7B Adapter mPLUG MiniGPT4 LLAVA-7B $0.0$ $-10.7$ 28.5+28.528.528.5 $-31.2$ $0.0$ 8.3+8.38.38.3 $-23.0$ $-5.8$ Adapter 30.9+30.930.930.9 $0.0$ 6.2+6.26.26.2 31.7+31.731.731.7 50.6+50.650.650.6 $0.0$ 6.8+6.86.86.8 $-14.3$ mPLUG-Owl 13.1+13.113.113.1 34.3+34.334.334.3 $0.0$ 6.6+6.66.66.6 1.5+1.51.51.5 $-18.4$ $0.0$ 13.5+13.513.513.5 MiniGPT-4 1.1+1.11.11.1 28.1+28.128.128.1 32.5+32.532.532.5 $0.0$ $-31.6$ 19.4+19.419.419.4 35.5+35.535.535.5 $0.0$
(b) Intersecting Lines & MathVista
Model Intersecting Lines MathVista LLAVA-7B Adapter mPLUG MiniGPT4 LLAVA-7B Adapter mPLUG MiniGPT4 LLAVA-7B $0.0$ 3.5+3.53.53.5 2.5+2.52.52.5 14.3+14.314.314.3 $0.0$ 12.2+12.212.212.2 2.4+2.42.42.4 $-7.7$ Adapter 2.3+2.32.32.3 $0.0$ 30.8+30.830.830.8 22.6+22.622.622.6 25.8+25.825.825.8 $0.0$ 3.3+3.33.33.3 20.0+20.020.020.0 mPLUG-Owl 7.3+7.37.37.3 $-12.6$ $0.0$ $-12.1$ 1.8+1.81.81.8 15.6+15.615.615.6 $0.0$ 16.0+16.016.016.0 MiniGPT-4 6.8+6.86.86.8 $-5.2$ $-9.7$ $0.0$ 2.5+2.52.52.5 6.0+6.06.06.0 3.7+3.73.73.7 $0.0$
Appendix D Prompts

### v2

@@ 6 Limitations
MTRE does have some limitations. First, it requires white-box access to the full sequence of early logits, so it cannot be applied when only final outputs or API-level confidences are available. Second, all experiments are conducted on four open-source VLMs and a handful of vision-question benchmarks; MTRE’s generality to other models, and other modalities (e.g. video), non-English prompts, or truly "in-the-wild" user queries remains untested.
@@ 7 Conclusion
Acknowledgment
This manuscript has been assigned LA-UR-25-30324. This research was funded by the Los Alamos National Laboratory (LANL) Laboratory Directed Research and Development (LDRD) program under grants 20230287ER, 20250850ECR and 20240868PRD3 and supported by LANL’s Institutional Computing Program, and by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001. We also thank Sambanova Systems for providing access to API calls for dataset synthesis utilized in this work
@@ Appendix A Datasets
We primarily evaluate or improvements on the datasets utilized by a first token linear probing technique discussed in Zhao et al. (2025). For each dataset, we construct a separate Type 2 dataset in the main text.
$M$-SafetyBench
MM-SafetyBench applies jailbreaking attacks to LVLMs across thirteen scenarios using malicious text prompts and images Liu et al. (2023b). The original dataset includes 1,680 unsafe questions for attacks, with each question generating three types of images: one created by Stable Diffusion Podell et al. (2023), one with rendered text, and one combining the first two. For our work, we use the augmented version of this dataset introduced in Zhao et al. (2025), which balances the dataset by adding a new data generation pipeline in MM-SafetyBench. This pipeline generates a total of 1,800 safe question-image pairs through GPT-4 prompts covering topics such as daily activities, economics, physical health, legal matters, politics, finance, sex, and government.
We train all models (MMD, Linear Probing, and MTRE) on these data to distinguish whether the output will be harmful. To remain consistent with Zhao et al. (2025), we also randomly select 10 samples from each category in both safe and unsafe sets and use 90 safe and 130 unsafe samples for training. The remaining data of the augmented MM-SafetyBench is used as the test set.
@@
MAD-Bench consists of 850 image-question pairs designed to deceive LVLMs. These deceptive pairs target various aspects, including object count, non-existent objects, object attributes, scene understanding, spatial relationships, and visual confusion Li et al. (2023). For example, given an image of two cats, a deceptive question might be: ’What are the three cats doing?’ In this case, rather than answering the question directly, the model should recognize the inconsistency between the question and the image. We also utilize an augmented dataset which adds an additional generated 1,000 normal questions based on the COCO val2017 dataset. We use 100 deceptive and 100 normal samples to train each proposed technique. The remaining data is then used as a validation dataset in each of our experiments.
@@
The MathVista dataset Lu et al. (2023) contains 1,000 image-question pairs related to math problems. This dataset challenges the model by requiring it to predict various types of answers, such as multiple-choice options, floating-point numbers, integers, and lists, making correctness prediction more complex. We prompt VLMs with the math visual prompts and evaluate their accuracy using GPT-4, following the scripts provided in the official GitHub repository.
Given the limited size of the dataset we implement a 4-fold cross-validation method to ensure the robustness of our analysis. In each fold, the model is provided with the output logits and trained to predict the accuracy of responses based on the logit distribution of each output token. Once trained, the model is applied to predict the accuracy of responses in the test segment. The performance of the model on this dataset is evaluated using the metrics discussed in Section 5.1 across all folds.
@@
Below we note the descriptions of the datasets given by Rahmanzadehgervi et al. (2024). Note that we alter each dataset primarily to experiment with more data, and more complicated cross-validation splits. We reduced the amount of shapes/diversity in all shape datasets due to the difficulty for smaller open-source models, and to reduce the mode collapse in VLM predictions. Similar to MathVista we implement a 4-fold cross-validation to account for the size of dataset. We are careful to not make each of the training splits identical to any of the validation splits for any of the folds.
  • Intersecting Lines: Following the work of Rahmanzadehgervi et al. (2024) we create 600 images of 2D line plots drawn on a white canvas. Each line plot consists of two line segments, defined by three points whose x-coordinates are fixed and equally spaced. The y-coordinates are randomly sampled to create two plots that intersect at exactly 0, 1 or 2 points. The goal of the VLM is to count the number of line intersections. There are 200 images with 0 intersections, 200 with 1 intersection, and 200 with 2 intersections. We denote explicit configurations in practice below:
@@
    • Shape Placement: The center of the nested stack is randomly positioned within the range [5,5][-5,5] for both $x$ and $y$ coordinates.
@@
A.1 Prompts